Let F(x,y,z)=⟨2xy,x^2+z,y⟩ be a conservative vector field. A potential function of F

is
Select one:
a. x^2y+yz+z


b. x^2yz


c. x^2y+yz−87

d. x^6yz


e. ⟨x^2y,yz,z⟩
1
Expert's answer
2020-06-08T18:47:37-0400

"F(x,y,z)=\u27e82xy,x^2+z,y\u27e9"

We know that "F" is the gradient of potential function "U(x,y,z)." So

"\dfrac{\partial U}{\partial x} = 2xy \Rightarrow U(x,y,z) = x^2y + f_1(y,z),
\\
\dfrac{\partial U}{\partial y} =x^2+z \Rightarrow U(x,y,z) = x^2y+yz+f_2(x,z),
\\
\dfrac{\partial U}{\partial z} = y \Rightarrow U(x,y,z) =yz+f_3(x,y)."

So "U(x,y,x) = x^2y+yz+\mathrm{const}." Therefore, we should choose answer c. "U(x,y,x) = x^2y+yz-87."